# Noncommutative weak type $(1,1)$ estimates for Calderón-Zygmund operators with $L_{2}$-mean Hörmander conditions

### v1

@@ 3. Two reductions
This section introduces two key lemmas (Lemma 3.1 and Lemma 3.2), which reduce the proof of the main theorem to the cases of real kernels and lacunary sequences. For their proofs, we refer the reader to [hong2023maximal].
@@
Let $T$ be a noncommutative Calderón-Zygmund operator (defined as in (1.8)) whose kernel $K$ satisfies (1.1) and the $L_{2}$-mean Hörmander condition (1.7). Let $T_{\epsilon}$ be defined as in (1.9). Then,
@@
Let $T$ be a noncommutative Calderón-Zygmund operator (defined as in (1.8)) whose kernel $K$ satisfies (1.1) and the $L_{2}$-mean Hörmander condition (1.7). Let $T_{\epsilon}^{i_{\epsilon}}$ be defined as in (3.2). Then
@@
The preceding two lemmas imply that Theorem 1.1 follows directly from the weak type (1,1) boundedness of $S_{i}$ given in (3.1). Thus, it remains to prove
@@
Let $T$ be a noncommutative Calderón-Zygmund operator defined as in (1.8) with the real kernel $K$ satisfying (1.1) and the $L_{2}$-mean Hörmander condition (1.7). Let $S_{i}$ be defined as in (3.1). Then, the sequence of operators $(S_{i})_{i\in\operatorname{\mathbb{Z}}}$ is of weak type $(1,1)$, that is, for any $f\in L_{1}(\operatorname{\mathcal{N}})$,
@@
4.2.1. The proof of Lemma 4.1
First, let us give an important lemma.
@@
Fix $1\leq q<\infty$ and $j\in\operatorname{\mathbb{Z}}$. Then, for $n\geq 1$,
$\sup_{y\in\operatorname{\mathbb{R}}^{d}}\frac{1}{|B_{r_{n-j}}|}\int_{B_{r_{n-j}}}\int_{\operatorname{\mathbb{R}}^{d}}$ $|K_{j}(x,y+z)-K_{j}(x,y)|^{q}{\rm d}x{\rm d}z$
$\lesssim 2^{-jd(q-1)}\big(H_{q}^{q}(n)+H_{q}^{q}(n+1)+2^{-qn}\big),$
where $r_{n-j}$ is defined in Theorem 2.4.
@@
For a fixed $z\in B_{r_{n-j}}$, we have
$\int_{\operatorname{\mathbb{R}}^{d}}|K_{j}(x,y+z)-K_{j}(x,y)|^{q}{\rm d}x$ $\lesssim\int_{\operatorname{\mathbb{R}}^{d}}|K(x,y+z)-K(x,y)|^{q}|\Phi_{j}(x-y)|^{q}{\rm d}x$
$\ \ +\int_{\operatorname{\mathbb{R}}^{d}}|K(x,y+z)|^{q}|\Phi_{j}(x-(y+z))-\Phi_{j}(x-y)|^{q}{\rm d}x$
$=:A_{1}+A_{2}.$
By the mean value theorem and $z\in B_{r_{n-j}}$,
$A_{2}\lesssim\int_{E_{j,n,y,z}}|K(x,y+z)|^{q}|z|^{q}2^{-jq}{\rm d}x\lesssim 2^{-qn}\int_{E_{j,n,y,z}}|K(x,y+z)|^{q}{\rm d}x,$
where $E_{j,n,y,z}$ denotes the support of the function $x\mapsto\big(\Phi_{j}(x-(y+z))-\Phi_{j}(x-y)\big)$. We claim that there exist two constants $C_{1,d}$, $C_{2,d}$ such that
(4.4) $E_{j,n,y,z}\subseteq\big\{x\in\operatorname{\mathbb{R}}^{d}:C_{1,d}2^{j}\leq|x-(y+z)|\leq C_{2,d}2^{j}\big\}.$
To prove this, note first that $\operatorname{supp}(\Phi_{j})\subset\big\{x\in\operatorname{\mathbb{R}}^{d}:2^{j-1}\sqrt{d}\leq|x|\leq 2^{j+1}\sqrt{d}\big\}$. Hence, when $\Phi_{j}(x-(y+z))-\Phi_{j}(x-y)\neq 0$, at least one of the following conditions holds:
@@ 2. $(ii)$
For case $(i)$, we have $|x-(y+z)|\approx 2^{j}$ immediately. For case $(ii)$, since $r_{n-j}\leq 2^{j-2}\sqrt{d}$ when $n\geq 1$,
$|x-(y+z)|\leq|x-y|+|z|\leq 2^{j+1}\sqrt{d}+2^{j-2}\sqrt{d},$
$|x-(y+z)|\geq|x-y|-|z|\geq 2^{j-1}\sqrt{d}-2^{j-2}\sqrt{d},$
which implyies $|x-(y+z)|\approx 2^{j}$. By the estimate for $E_{j,n,y,z}$ as above with the size condition (1.1), we obtain a bound on $A_{2}$:
$A_{2}\lesssim 2^{-jd(q-1)}\cdot 2^{-qn}.$
For $A_{1}$, the definition of $r_{n-j}$ gives
$A_{1}$ $\leq\int_{2^{j-1}\sqrt{d}\leq|x-y|\leq 2^{j+1}\sqrt{d}}|K(x,y+z)-K(x,y)|^{q}{\rm d}x$
$=\int_{2^{n}r_{n-j}\leq|x-y|\leq 2^{n+1}r_{n-j}}|K(x,y+z)-K(x,y)|^{q}{\rm d}x$
$\quad+\int_{2^{n+1}r_{n-j}\leq|x-y|\leq 2^{n+2}r_{n-j}}|K(x,y+z)-K(x,y)|^{q}{\rm d}x.$
Finally, combining the estimates for $A_{1}$, $A_{2}$, and the $L_{2}$-mean Hörmander condition (1.7), we get
$\frac{1}{|B_{r_{n-j}}|}\int_{B_{r_{n-j}}}\int_{\operatorname{\mathbb{R}}^{d}}|K_{j}(x,y+z)-K_{j}(x,y)|^{q}{\rm d}x{\rm d}z$
$\quad\leq\frac{1}{|B_{r_{n-j}}|}\int_{B_{r_{n-j}}}\Big(\int_{2^{n}r_{n-j}\leq|x-y|\leq 2^{n+1}r_{n-j}}|K(x,y+z)-K(x,y)|^{q}{\rm d}x\Big){\rm d}z$
$\quad\quad+\frac{1}{|B_{r_{n-j}}|}\int_{B_{r_{n-j}}}\Big(\int_{2^{n+1}r_{n-j}\leq|x-y|\leq 2^{n+2}r_{n-j}}|K(x,y+z)-K(x,y)|^{q}{\rm d}x\Big){\rm d}z$
$\quad\quad+2^{-jd(q-1)}\cdot 2^{-qn}$
$\quad\lesssim 2^{-jd(q-1)}\big(H_{q}^{q}(n)+H_{q}^{q}(n+1)+2^{-qn}\big),$
which completes the proof. ∎
Proof of Lemma 4.1.
@@
Indeed, for a fixed $x\in\operatorname{\mathbb{R}}^{d}$, Remark 2.7 ($s=[6\sqrt{d}]$) and Lemma 2.6 $(ii)$ yield
$\zeta(x)\mathcal{T}_{j}\widetilde{b}_{\text{d},n-j}(x)\zeta(x)=\int_{\operatorname{\mathbb{R}}^{d}}K_{j}(x,y)\Big(\zeta(x)\widetilde{b}_{\text{d},n-j}(y)\zeta(x)\Big)\chi_{\big\{|x-y|>\big(\frac{2[6\sqrt{d}]+1}{\sqrt{d}}-2\big)r_{n-j}\big\}}(y){\rm d}y.$
For $n\leq 1$, the definition of $r_{n-j}$ (see Theorem 2.4) directly implies
$2^{j+1}\sqrt{d}=2^{n+2}r_{n-j}<\Big(\frac{2[6\sqrt{d}]+1}{\sqrt{d}}-2\Big)r_{n-j}.$
@@
which shows that $\zeta\mathcal{T}_{j}\widetilde{b}_{\text{d},n-j}\zeta=0$. A similar argument works for $\zeta\mathcal{T}_{j}b_{\text{d},n-j}^{0}\zeta$. Thus, we obtain (4.5)
@@
$=\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\int_{\operatorname{\mathbb{R}}^{d}}\int_{\operatorname{\mathbb{R}}^{d}}\big(K_{j}(x,y)-K_{j}(x,c(Q)\big)b_{\text{d},n-j}^{Q}(u)h_{r_{n-j}}(y-u){\rm d}y{\rm d}u.$
By the facts that $\operatorname{supp}(h_{r_{n-j}})\subseteq B_{r_{n-j}}$, $\operatorname{supp}(b^{Q}_{\text{d},n-j})\subseteq Q$, and the Minkowski inquality, we have
@@
$\lesssim\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\int_{\operatorname{\mathbb{R}}^{d}}\frac{\left\|h\right\|_{\infty}}{(r_{n-j})^{d}}\Big(\int_{Q}\int_{B_{r_{n-j}}(u)}|K_{j}(x,y)-K_{j}(x,c(Q))|\|b_{\text{d},n-j}^{Q}(u)\|_{L_{1}(\operatorname{\mathcal{M}})}{\rm d}y{\rm d}u\Big){\rm d}x$
$\lesssim\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\int_{Q}\|b_{\text{d},n-j}^{Q}(u)\|_{L_{1}(\operatorname{\mathcal{M}})}\Big(\frac{1}{|B_{r_{n-j}}|}\int_{B_{r_{n-j}}(u)}\int_{\operatorname{\mathbb{R}}^{d}}|K_{j}(x,y)-K_{j}(x,c(Q))|{\rm d}x{\rm d}y\Big){\rm d}u.$
For a fixed cube $Q\in\operatorname{\mathcal{Q}}_{n-j}$, note that the variable $u$ in the outer integral is restricted to $Q$. Moverover, by the definition of $r_{n-j}$,
$|y-c(Q)|\leq|y-u|+|u-c(Q)|\leq r_{n-j}+r_{n-j}=r_{n-j-1}.$
This implies that whenever $u\in Q$ and $y\in B_{r_{n-j}}(u)$, we have $y\in B_{r_{n-j-1}}(c(Q))$. Therefore, let z=yc(Q)z=y-c(Q),
$\frac{1}{|B_{r_{n-j}}|}$ $\int_{B_{r_{n-j}}(u)}\int_{\operatorname{\mathbb{R}}^{d}}|K_{j}(x,y)-K_{j}(x,c(Q))|{\rm d}x{\rm d}y$
(4.6) $\ \lesssim\frac{1}{|B_{r_{n-j-1}}|}\int_{B_{r_{n-j-1}}}\int_{\operatorname{\mathbb{R}}^{d}}|K_{j}(x,z+c(Q))-K_{j}(x,c(Q))|{\rm d}x{\rm d}z.$
By employing Lemma 4.3 with $y=c(Q)$, $q=1$, and setting $n$ to n1n-1, (4.2.1) is bounded by
$H_{1}(n-1)+H_{1}(n)+2^{-n+1}.$
Then, applying the inequality $H_{1}(n)\lesssim_{d}H_{2}(n)$, the $L_{2}$-mean Hörmander condition (1.7) and the boundedness of $b_{\text{d},n}^{Q}$ in Lemma 2.6 $(ii)$, we obtain
$\ \sum_{n\geq 2}\sum_{j\leq n-1}\big\|\zeta\mathcal{T}_{j}\widetilde{b}_{\text{d}}\zeta\big\|_{L_{1}(\operatorname{\mathcal{N}})}$
$\ \lesssim\sum_{n\geq 2}\sum_{j\leq n-1}\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\int_{Q}\big\|b_{\text{d},n-j}^{Q}(u)\big\|_{L_{1}(\operatorname{\mathcal{M}})}\left(H_{1}(n-1)+H_{1}(n)+2^{-n+1}\right){\rm d}u$
$\ \lesssim\sum_{n\geq 1}\big(2H_{2}(n)+2^{-n+1}\big)\sum_{m\geq 1}\sum_{Q\in\operatorname{\mathcal{Q}}_{m}}\int_{Q}\big\|b_{\text{d},m}^{Q}(u)\big\|_{L_{1}(\operatorname{\mathcal{M}})}{\rm d}u$
@@
$(ii)$ Again by Remark 2.8 and (4.5), it remains to prove:
@@
Fix a cube $Q\in\operatorname{\mathcal{Q}}_{n-j}$ and recall the definition of $\widetilde{b}_{\text{d},n-j}^{Q}$ in Theorem 2.4,
$\mathcal{T}_{j}\widetilde{b}_{\text{d},n-j}^{Q}(x)$ $=\int_{\operatorname{\mathbb{R}}^{d}}K_{j}(x,y)\int_{\operatorname{\mathbb{R}}^{d}}b_{\text{d},n-j}^{Q}(y-z)h_{r_{n-j}}(z){\rm d}z{\rm d}y$
$=\int_{\operatorname{\mathbb{R}}^{d}}h_{r_{n-j}}(z)\int_{\operatorname{\mathbb{R}}^{d}}K_{j}(x,y+z)b_{\text{d},n-j}^{Q}(y){\rm d}y{\rm d}z.$
Moreover, by $\int_{\operatorname{\mathbb{R}}^{d}}h=1$,
$\mathcal{T}_{j}b_{\text{d},n-j}^{Q}(x)=\int_{\operatorname{\mathbb{R}}^{d}}h_{r_{n-j}}(z)\int_{\operatorname{\mathbb{R}}^{d}}K_{j}(x,y)b_{\text{d},n-j}^{Q}(y){\rm d}y{\rm d}z.$
Noting again the supports of $h_{r_{n-j}}$, $b_{\text{d},n-j}^{Q}$ and using the Minkowski inequality, we get
@@
$\ \lesssim\int_{\operatorname{\mathbb{R}}^{d}}\Big(\frac{1}{|B_{r_{n-j}}|}\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\int_{Q}\int_{B_{r_{n-j}}}|K_{j}(x,y+z)-K_{j}(x,y)|\|b_{\text{d},n-j}^{Q}(y)\|_{L_{1}(\operatorname{\mathcal{M}})}{\rm d}z{\rm d}y\Big){\rm d}x$
$\ \lesssim\big(H_{1}(n)+H_{1}(n+1)+2^{-n}\big)\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\|b_{\text{d},n-j}^{Q}\|_{L_{1}(\operatorname{\mathcal{N}})},$
where the last inequality follows from Lemma 4.3 with $q=1$. Finally, an application of the fact that $H_{1}(n)\lesssim_{d}H_{2}(n)$, together with the $L_{2}$-mean Hörmander condition (1.7), and the boundedness of $b_{\mathrm{d},n}^{Q}$ in Lemma 2.6 $(ii)$, yield
$\sum_{n\geq 2}\sum_{j\leq n-1}\left\|\zeta\mathcal{T}_{j}b_{\text{d},n-j}^{0}\zeta\right\|_{L_{1}(\operatorname{\mathcal{N}})}$
$\quad\leq\sum_{n\geq 2}\sum_{j\leq n-1}\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\|b_{\text{d},n-j}^{Q}\|_{L_{1}(\operatorname{\mathcal{N}})}\left(H_{1}(n)+H_{1}(n+1)+2^{-n}\right)$
$\quad\lesssim\sum_{n\geq 2}\left(2H_{2}(n)+2^{-n}\right)\sum_{m\geq 1}\sum_{Q\in\operatorname{\mathcal{Q}}_{m}}\|b_{\text{d},m}^{Q}\|_{L_{1}(\operatorname{\mathcal{N}})}$
$\quad\lesssim\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})},$
which completes the argument. ∎

### v2

@@ 3. Two reductions
This section introduces two key lemmas (Lemma 3.1 and Lemma 3.2), which reduce the proof of the main theorem to the cases of real kernels and lacunary sequences. For their proofs, we refer the reader to [7].
@@
Let $T$ be a noncommutative Calderón-Zygmund operator (defined as in (1.7)) whose kernel $K$ satisfies (1.1) and the $L_{1}$-integral condition (1.6). Let $T_{\epsilon}$ be defined as in (1.8). Then,
@@
Let $T$ be a noncommutative Calderón-Zygmund operator (defined as in (1.7)) whose kernel $K$ satisfies (1.1) and the $L_{1}$-integral condition (1.6). Let $T_{\epsilon}^{i_{\epsilon}}$ be defined as in (3.3). Then
@@
The preceding two lemmas imply that Theorem 1.1 follows directly from the weak type (1,1) boundedness of $S_{i}$ given in (3.2). Thus, it is enough to prove
@@
Let $T$ be a noncommutative Calderón-Zygmund operator defined as in (1.7) with the real kernel $K$ satisfying (1.1) and the $L_{1}$-integral condition (1.6). Let $S_{i}$ be defined as in (3.2). Then, the sequence of operators $(S_{i})_{i\in\operatorname{\mathbb{Z}}}$ is of weak type $(1,1)$, that is, for any $f\in L_{1}(\operatorname{\mathcal{N}})$,
@@
We first introduce a lemma that will be essential for proving Lemma 4.1 and Lemma 4.2. For a fixed dyadic cube $Q$, we recall the kernel average
$\mathcal{K}_{Q}(x,y)=\frac{1}{|Q|}\int_{|v|\leq\frac{\sqrt{d}}{2}\ell(Q)}|K(x,y+v)-K(x,y)|{\rm d}v.$
When there is no risk of confusion, it is convenient to denote by $\mathcal{K}_{j,Q}$ the corresponding average for the truncated kernel $K_{j}$:
$\mathcal{K}_{j,Q}(x,y):=\frac{1}{|Q|}\int_{|z|\leq\frac{\sqrt{d}}{2}\ell(Q)}|K_{j}(x,y+z)-K_{j}(x,y)|{\rm d}z.$
@@
Fix a dyadic cube $Q\in\operatorname{\mathcal{Q}}_{n-j}$ and $j\in\operatorname{\mathbb{Z}}$. Then, for $n\geq 3$,
$\int_{\operatorname{\mathbb{R}}^{d}}\sup_{y\in Q}\mathcal{K}_{j,Q}(x,y){\rm d}x\lesssim\mathcal{H}(n-2)+\mathcal{H}(n-1)+\mathcal{H}(n)+\mathcal{H}(n+1)+2^{-n}.$
@@
From the definition of $K_{j}$ in (3.1), we have
$\mathcal{K}_{j,Q}(x,y)$ $\leq\frac{1}{|Q|}\Big(\int_{|z|\leq\frac{\sqrt{d}}{2}\ell(Q)}|K(x,y+z)-K(x,y)||\Phi_{j}(x-y)|{\rm d}z$
$\quad+\int_{|z|\leq\frac{\sqrt{d}}{2}\ell(Q)}|K(x,y+z)||\Phi_{j}(x-(y+z))-\Phi_{j}(x-y)|{\rm d}z\Big).$
Consequently,
$\int_{\operatorname{\mathbb{R}}^{d}}\sup_{y\in Q}\mathcal{K}_{j,Q}(x,y){\rm d}x\leq A_{1}+A_{2},$
where
$A_{1}$ $:=\frac{1}{|Q|}\int_{\operatorname{\mathbb{R}}^{d}}\sup_{y\in Q}\Big(\int_{|z|\leq\frac{\sqrt{d}}{2}\ell(Q)}|K(x,y+z)-K(x,y)|{\rm d}z|\Phi_{j}(x-y)|\Big){\rm d}x$
$=\int_{\operatorname{\mathbb{R}}^{d}}\sup_{y\in Q}\Big(\mathcal{K}_{Q}(x,y)|\Phi_{j}(x-y)|\Big){\rm d}x,$
and
$A_{2}:=\frac{1}{|Q|}\int_{\operatorname{\mathbb{R}}^{d}}\Big(\sup_{y\in Q}\int_{|z|\leq\frac{\sqrt{d}}{2}\ell(Q)}|K(x,y+z)||\Phi_{j}(x-(y+z))-\Phi_{j}(x-y)|{\rm d}z\Big){\rm d}x.$
We first estimate $A_{1}$. When $y\in Q$, by $\operatorname{supp}(\Phi_{j})\subset\big\{x\in\operatorname{\mathbb{R}}^{d}:2^{j-1}\sqrt{d}\leq|x|\leq 2^{j+1}\sqrt{d}\big\}$ and
(4.4) $\frac{\sqrt{d}}{2}\ell(Q)=2^{-n+j-1}\sqrt{d}\leq 2^{j-4}\sqrt{d},\quad\text{when}\ n\geq 3,$
we obtain
$|x-c(Q)|\leq|x-y|+|y-c(Q)|\leq 2^{j+1}\sqrt{d}+(\sqrt{d}/2)\ell(Q)\leq 2^{j+2}\sqrt{d},$
$|x-c(Q)|\geq|x-y|-|y-c(Q)|\leq 2^{j-1}\sqrt{d}-(\sqrt{d}/2)\ell(Q)\leq 2^{j-2}\sqrt{d}.$
This means,
$2^{n-2}\sqrt{d}\ell(Q)\leq|x-c(Q)|\leq 2^{n+2}\sqrt{d}\ell(Q).$
Hence,
$A_{1}$ $\leq\int_{2^{n-2}\sqrt{d}\ell(Q)\leq|x-c(Q)|\leq 2^{n+2}\sqrt{d}\ell(Q)}\sup_{y\in Q}\mathcal{K}(x,y){\rm d}x$
$=\sum_{k=-2}^{1}\int_{2^{n+k}\sqrt{d}\ell(Q)\leq|x-c(Q)|\leq 2^{n+k+1}\sqrt{d}\ell(Q)}\sup_{y\in Q}\mathcal{K}(x,y){\rm d}x$
$\leq\mathcal{H}(n-2)+\mathcal{H}(n-1)+\mathcal{H}(n)+\mathcal{H}(n+1).$
Consider $A_{2}$. We first claim that there exists a constant $C_{d}$ such that the function $x\mapsto\big(\Phi_{j}(x-(y+z))-\Phi_{j}(x-y)\big)$ is supported in the set
(4.5) $\big\{x\in\operatorname{\mathbb{R}}^{d}:|x-c(Q)|\leq C_{d}2^{j}\big\}=:E_{j,Q},$
whenever $y\in Q$ and $z\in B_{\frac{\sqrt{d}}{2}\ell(Q)}$. To prove this, by the support of $\Phi_{j}$, when $\Phi_{j}(x-(y+z))-\Phi_{j}(x-y)\neq 0$, at least one of the following conditions holds:
@@ 2. $(ii)$
In case $(i)$, it follows from $|z|\leq\frac{\sqrt{d}}{2}\ell(Q)$, $y\in Q$ and (4.4) that
$|x-c(Q)|\leq|x-(y+z)|+|z|+|y-c(Q)|\leq 2^{j+1}\sqrt{d}+2^{j-4}\sqrt{d}+2^{j-4}\sqrt{d}.$
In case $(ii)$, similarly,
$|x-c(Q)|\leq|x-y|+|y-c(Q)|\leq 2^{j+1}\sqrt{d}+2^{j-4}\sqrt{d}.$
Thus (4.5) holds.
Moreover, when $\Phi_{j}(x-(y+z))-\Phi_{j}(x-y)\neq 0$, we have another useful estimate:
(4.6) $|x-(y+z)|\approx 2^{j}.$
Indeed, this is immediate in case $(i)$. And in case $(ii)$, using $|z|\leq\frac{\sqrt{d}}{2}\ell(Q)$ and (4.4) again, we have
$|x-(y+z)|\leq|x-y|+|z|\leq 2^{j+1}\sqrt{d}+2^{j-4}\sqrt{d},$
$|x-(y+z)|\geq|x-y|-|z|\geq 2^{j-1}\sqrt{d}-2^{j-4}\sqrt{d},$
which also implies $|x-(y+z)|\approx 2^{j}$, leading to (4.6).
Now, we can bound $A_{2}$. By the size condition (1.1), the mean-value theorem, the bound $|z|\leq\frac{\sqrt{d}}{2}\ell(Q)=2^{-n+j-1}\sqrt{d}$, together with (4.5) and (4.6),
$A_{2}$ $\lesssim\int_{E_{j,Q}}\Big(\sup_{y\in Q}\frac{1}{|Q|}\int_{|z|\leq\frac{\sqrt{d}}{2}\ell(Q)}\frac{|z|2^{-j}}{|x-(y+z)|^{d}}{\rm d}z\Big){\rm d}x$
$\lesssim 2^{-n-jd}\int_{E_{j,Q}}\Big(\sup_{y\in Q}\frac{1}{|Q|}\int_{|z|\leq\frac{\sqrt{d}}{2}\ell(Q)}1{\rm d}z\Big){\rm d}x$
$\lesssim 2^{-n}.$
This completes the proof. ∎
4.2.1. The proof of Lemma 4.1
@@
Indeed, for a fixed $x\in\operatorname{\mathbb{R}}^{d}$, Lemma 2.5 $(ii)$ and Remark 2.6 ($s=[50\sqrt{d}]$) implies
$\zeta(x)\mathcal{T}_{j}\widetilde{b}_{\text{d},n-j}(x)\zeta(x)$
$\quad=\int_{\operatorname{\mathbb{R}}^{d}}K_{j}(x,y)\Big(\zeta(x)\widetilde{b}_{\text{d},n-j}(y)\zeta(x)\Big)\chi_{\big\{|x-y|>([50\sqrt{d}]-\sqrt{d})2^{-n+j}\big\}}(y){\rm d}y.$
@@
when
$|x-y|\geq([50\sqrt{d}]-\sqrt{d})2^{-n+j}\geq([50\sqrt{d}]-\sqrt{d})2^{j-3}\geq 2^{j+1}\sqrt{d},$
where the second inequality uses $n\leq 3$. This shows that $\zeta\mathcal{T}_{j}\widetilde{b}_{\text{d},n-j}\zeta=0$ when $n\leq 3$. The same argument works for $\zeta\mathcal{T}_{j}b_{\text{d},n-j}^{0}\zeta$, giving rise to (4.7).
@@
$=\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\int_{\operatorname{\mathbb{R}}^{d}}\int_{\operatorname{\mathbb{R}}^{d}}\big(K_{j}(x,y)-K_{j}(x,c(Q)\big)b_{\text{d},n-j}^{Q}(u)h_{\frac{\sqrt{d}}{2}\ell(Q)}(y-u){\rm d}y{\rm d}u.$
Then, by $\operatorname{supp}(h_{\frac{\sqrt{d}}{2}\ell(Q)})\subseteq B_{\frac{\sqrt{d}}{2}\ell(Q)}$, $\operatorname{supp}(b^{Q}_{\text{d},n-j})\subseteq Q$, and using the Minkowski inequality,
@@
$\lesssim\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\int_{\operatorname{\mathbb{R}}^{d}}\Big(\frac{\left\|h\right\|_{\infty}}{\ell(Q)^{d}}\int_{Q}\int_{B_{\frac{\sqrt{d}}{2}\ell(Q)}(u)}|K_{j}(x,y)-K_{j}(x,c(Q))|\|b_{\text{d},n-j}^{Q}(u)\|_{L_{1}(\operatorname{\mathcal{M}})}{\rm d}y{\rm d}u\Big){\rm d}x$
$\lesssim\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\int_{Q}\Big(\int_{\operatorname{\mathbb{R}}^{d}}\frac{1}{|Q|}\int_{B_{\frac{\sqrt{d}}{2}\ell(Q)}(u)}|K_{j}(x,y)-K_{j}(x,c(Q))|{\rm d}y{\rm d}x\Big)\|b_{\text{d},n-j}^{Q}(u)\|_{L_{1}(\operatorname{\mathcal{M}})}{\rm d}u.$
Fix a cube $Q\in\operatorname{\mathcal{Q}}_{n-j}$. For $y\in B_{\frac{\sqrt{d}}{2}\ell(Q)}(u)$, we have
$|y-c(Q)|\leq|y-u|+|u-c(Q)|\leq\frac{\sqrt{d}}{2}\ell(Q)+\frac{\sqrt{d}}{2}\ell(Q)\leq\sqrt{d}\ell(Q)$
whenever $u\in Q$. Hence, by letting z=yc(Q)z=y-c(Q),
$\frac{1}{|Q|}\int_{B_{\frac{\sqrt{d}}{2}\ell(Q)}(u)}|K_{j}(x,y)-K_{j}(x,c(Q))|{\rm d}y$
(4.8) $\quad\lesssim\frac{1}{|Q|}\int_{|z|\leq\sqrt{d}\ell(Q)}|K_{j}(x,z+c(Q))-K_{j}(x,c(Q))|{\rm d}z.$
Now, let the cube $\widehat{Q}$ be the unique father cube of $Q$ in $\operatorname{\mathcal{Q}}_{n-j-1}$. This implies
[…]
$\leq\sup_{y\in\widehat{Q}}\mathcal{K}_{j,\widehat{Q}}(x,y),$
which leads to
$\int_{\operatorname{\mathbb{R}}^{d}}\frac{1}{|Q|}\int_{B_{\frac{\sqrt{d}}{2}\ell(Q)}(u)}|K_{j}(x,y)-K_{j}(x,c(Q))|{\rm d}y{\rm d}x$
$\quad\lesssim\int_{\operatorname{\mathbb{R}}^{d}}\sup_{y\in\widehat{Q}}\mathcal{K}_{j,\widehat{Q}}(x,y){\rm d}x$
$\quad\lesssim\mathcal{H}(n-3)+\mathcal{H}(n-2)+\mathcal{H}(n-1)+\mathcal{H}(n)+2^{-n+1},$
where the last bound comes from Lemma 4.3 with $n$ replaced by n1n-1. Finally, applying the $L_{1}$-integral condition (1.6) and the boundedness of $b_{\text{d},n}^{Q}$ in Lemma 2.5 $(ii)$, we obtain
$\sum_{n\geq 4}\sum_{j\leq n-1}\big\|\zeta\mathcal{T}_{j}\widetilde{b}_{\text{d}}\zeta\big\|_{L_{1}(\operatorname{\mathcal{N}})}$ $\lesssim\sum_{n\geq 4}\sum_{j\leq n-1}\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\int_{Q}\big\|b_{\text{d},n-j}^{Q}(u)\big\|_{L_{1}(\operatorname{\mathcal{M}})}{\rm d}u$
$\quad\quad\left(\mathcal{H}(n-3)+\mathcal{H}(n-2)+\mathcal{H}(n-1)+\mathcal{H}(n)+2^{-n+1}\right)$
$\ \lesssim\sum_{n\geq 4}\big(4\mathcal{H}(n-3)+2^{-n+1}\big)\sum_{m\geq 1}\sum_{Q\in\operatorname{\mathcal{Q}}_{m}}\big\|b_{\text{d},m}^{Q}\big\|_{L_{1}(\operatorname{\mathcal{N}})}$
@@
$(ii)$ Again by Remark 2.7 and (4.7), it suffices to prove:
@@
Fix a cube $Q\in\operatorname{\mathcal{Q}}_{n-j}$. From the definition of $\widetilde{b}_{\text{d},n-j}^{Q}$ in Theorem 2.4,
$\mathcal{T}_{j}\widetilde{b}_{\text{d},n-j}^{Q}(x)$ $=\int_{\operatorname{\mathbb{R}}^{d}}K_{j}(x,y)\Big(\int_{\operatorname{\mathbb{R}}^{d}}b_{\text{d},n-j}^{Q}(y-z)h_{\frac{\sqrt{d}}{2}\ell(Q)}(z){\rm d}z\Big){\rm d}y$
$=\int_{\operatorname{\mathbb{R}}^{d}}\Big(\int_{\operatorname{\mathbb{R}}^{d}}K_{j}(x,y+z)h_{\frac{\sqrt{d}}{2}\ell(Q)}(z){\rm d}z\Big)b_{\text{d},n-j}^{Q}(y){\rm d}y.$
Moreover, by $\int_{\operatorname{\mathbb{R}}^{d}}h_{\frac{\sqrt{d}}{2}\ell(Q)}=1$,
$\mathcal{T}_{j}b_{\text{d},n-j}^{Q}(x)=\int_{\operatorname{\mathbb{R}}^{d}}\Big(\int_{\operatorname{\mathbb{R}}^{d}}K_{j}(x,y)h_{\frac{\sqrt{d}}{2}\ell(Q)}(z){\rm d}z\Big)b_{\text{d},n-j}^{Q}(y){\rm d}y.$
Noting again the supports of $h_{\frac{\sqrt{d}}{2}\ell(Q)}$, $b_{\text{d},n-j}^{Q}$ and using the Minkowski inequality, we get
@@
$\ \lesssim\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\int_{Q}\int_{\operatorname{\mathbb{R}}^{d}}\Big(\frac{1}{|Q|}\int_{|z|\leq\frac{\sqrt{d}}{2}\ell(Q)}|K_{j}(x,y+z)-K_{j}(x,y)|{\rm d}z\Big){\rm d}x\|b_{\text{d},n-j}^{Q}(y)\|_{L_{1}(\operatorname{\mathcal{M}})}{\rm d}y$
$\ \lesssim\big(\mathcal{H}(n-2)+\mathcal{H}(n-1)+\mathcal{H}(n)+\mathcal{H}(n+1)+2^{-n}\big)\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\|b_{\text{d},n-j}^{Q}\|_{L_{1}(\operatorname{\mathcal{N}})},$
where the last inequality follows from Lemma 4.3. Finally, an application of the $L_{1}$-integral condition (1.6), and the boundedness of $b_{\mathrm{d},n}^{Q}$ in Lemma 2.5 $(ii)$, yield
$\sum_{n\geq 4}\sum_{j\leq n-1}\left\|\zeta\mathcal{T}_{j}b_{\text{d},n-j}^{0}\zeta\right\|_{L_{1}(\operatorname{\mathcal{N}})}$ $\lesssim\sum_{n\geq 4}\left(4\mathcal{H}(n-2)+2^{-n}\right)\sum_{m\geq 1}\sum_{Q\in\operatorname{\mathcal{Q}}_{m}}\|b_{\text{d},m}^{Q}\|_{L_{1}(\operatorname{\mathcal{N}})}$
$\lesssim\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})},$
which completes the argument.